OpenStudy (anonymous):

Mike bought a shirt and a package of socks. The short cost 17.50 more than the socks. Mike only spent 30 dollars. How do i set this up. I get the x+17.50

OpenStudy (anonymous):

I assume you want to find the value of the socks so you set it up like:\[x + 17.50 = 30\]

OpenStudy (anonymous):

Then do you know how to find x?

OpenStudy (anonymous):

hey mwrnnj are u good at lit. english

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Yes i understand how to find x but it also wants me to find the value of the shirt. so i dont know what to do once i find x

OpenStudy (anonymous):

oh ok one sec

OpenStudy (anonymous):

x is the socks, (x+17.50) is the shirt, set = to 30 and solve. x+(x+17.50)=30

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

Sorry I didn't see that, let me check. Thank you though.

OpenStudy (anonymous):

YES THANK YOU !

OpenStudy (anonymous):

:)